40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Jun 20 16:36:44 2014, Nichin, Update, Electronics, RF cables removed 
    Reply  Tue Jun 24 16:52:43 2014, Nichin, Update, Electronics, An RF cable re-installed 
Message ID: 10093     Entry time: Tue Jun 24 16:52:43 2014     In reply to: 10082
Author: Nichin 
Type: Update 
Category: Electronics 
Subject: An RF cable re-installed 

Quote:

 [Nichin, Eric G]

As mentioned in Elog 10062, we found RF cables running between demodulators in rack 1Y2 and RF switch in 1Y1 to have bad SMA connectors (No shield / bad soldering / no caps).

we pulled out all the cables belonging to PD frequency response measurement system , 8 in total, and all of them about 5.5m in length.

Their labels read :

REFL33, REFL11, REFL55, AS55, POX11, REFL165, POP22 and POP110. 

All of them are now sitting inside a plastic box in the contorl room.

On another note, instead of fixing all the cables ourselves, Steve and Eric G decided to order custom made RF cables from Pasternack as professionally soldered cables are worth it. We have placed an order for 2 cables (RG405-550CM) to check out  and test them before we order all of the cables.

 The new RF cables arrived. But unfortunately we did not realize that RG405 was a Semi-rigid coax cable, with a copper shielding. These are meant to be installed in setups that will not be changed / disturbed. We need to order a different set of cables. The new cables have joined the other cables in the plastic box mentioned above.

For now to check if the old setup is still working, I have installed an RF cable (that we earlier pulled out and looks like in good shape, labelled REFL33) between the AS55 Demodulator output PD RF MON in rack 1Y2 and the network analyzer input. Since Manasa and the others were busy working with the interferometer, I did not switch on the laser and did not take any readings. The power supply to REF DET remains off.

I will continue with the measurements tomorrow morning and also try to get the data wirelessly using Alex's code. 

ELOG V3.1.3-